(2 HarKS) QUESTION 3 (4 marks) (c) You have been offered the chance to invest in a film venture. You have to put up $500,000 in capital to earn $300,000 in year one and $300,000 in year two. What rate of interest will you earn if you take this opportunity? (Show all working) (2 marks) (d) A company borrowed money by selling promissory notes with 90 days to maturity. If each note has a face value S100,000 and was sold for $98,303, what was the current market yield on promissory notes in this risk class at issuance? (2 marks) Distributing prohibited | Downloaded by Dominic Ho (weijian73@gmail.com) BU2004/BX2014:03/2/15
Answer #1

(c).

We have following information -

Investment - $ 500,000

1st year cash inflow - $ 300,000

2nd year cash inflow - $ 300,000

Here, interest rate is the rate at which Present value of Future cash inflows equals to the Initial Investment.

Suppose interest rate is i,

then,

300, 000 300,000 500,000 1

By solving this we will get interest rate.

Alternatively, we can also calculate the interest rate in excel with "=irr" formula. Please refer below spreadsheet for calculation -

C6 Year C) 1 2 Cash flows $ (500,000) $300,000 $300,000 13.07% 4 Interest rate

Formula reference -

C6 fIRR(C3:C5) Year Cash flows 0 1 2 Interest rate 500000 300000 300000 4 -IRR(C3:C5)

Thus, Interest rate is 13.07%

(d)

Present Value of Maturity amount of promissory note at yield rate is equal to current sale price of promissory note.

Suppose yield rate is Y for 90 days

then,

Facevalue SaIeprice of Note = (1 + Y))

98,303 = rac{100,000}{(1+Y)^1}

(1+Y) = 100,000/98,303

(1+Y) = 1.01726

Y = 1.01726-1

Y = 0.01726

Thus, Annual Yield rate would be = (0.01726/90)*365

= 0.06999

Yield rate p.a = 7%

(rounded off)

Please note- we can also calculate yield rate in excel using "=irr " formula as calculated in question-(c)